OpenStudy (anonymous):

What is the solution set for the following quadratic equation? x2 - 4x + 4 = 0 {-2, 2} {-2} {1, 3} {2}

OpenStudy (anonymous):

2

OpenStudy (anonymous):

2 ^2 is squared (+-) is like plus or minus... sorry there's no key for that :) (-b(+-)√b^2-4AC)/2A a=1 b=-4 c=4 (4(+-)√4^2-4(1)(4))/2(1) (4(+-)√16-16)/2 (4(+-)√0)/2 (4(+-)0)/2 since 4 (+-) 0 is always 4, we can cancel that out 4/2 2

OpenStudy (whpalmer4):

\[x^2 - 4x + 4 = 0\]We can factor that if we can think of a pair of factors of 4 that add to -4. Obviously, both factors must be negative to have a product of +4. -2 and -2 is such a pair: -2 + -2 = -4, and -2 * -2 = +4. \[x^2-4x+4 = (x-2)(x-2)\] \[(x-2)(x-2) = 0\] For that to be true, at least one of the product terms must = 0: \[x-2=0\]\[x-2+2 = 0+2\]\[x=2\] Checking our answer: \[x^2-4x+4=0\]\[(2)^2-4(2)+4=0\]\[4-8+4=0\]\[0=0\checkmark\] Therefore the solution set is {2}